A ball falls freely from a height 1 m on the ground and rebounces to a height of0.8 .find the coefficient of restitution

Grade:11

1 Answers

Pranav Jayant Chiddarwar
6 years ago
velocity of approach = root(2gh) = root(2*10*1) =  root(20)
velocity of seperation =root(2gh)b = root(2*10*0.8) = root(16)
e = velocity of approach/velocity of seperation
e =  root(20)/root(16) = [root(5)]/2 = 1.1